# Safe Text-to-Image Generation: Simply Sanitize the Prompt Embedding

### v1

@@ Abstract
Given that semantic consistency between input text and output image is a fundamental requirement for T2I models, we identify that textual representations (i.e., prompt embeddings) are likely the primary source of unsafe generation. To this end, we propose a vision-agnostic safe generation framework, Embedding Sanitizer (ES), which focuses on erasing inappropriate concepts from prompt embeddings and uses the sanitized embeddings to guide the model for safe generation. ES is applied to the output of the text encoder as a plug-and-play module, enabling seamless integration with different T2I models as well as other safeguards. In addition, ES’s unique scoring mechanism assigns a score to each token in the prompt to indicate its potential harmfulness, and dynamically adjusts the sanitization intensity to balance defensive performance and generation quality. Through extensive evaluation on five prompt benchmarks, our approach achieves state-of-the-art robustness by sanitizing the source (prompt embedding) of unsafe generation compared to nine baseline methods. It significantly outperforms existing safeguards in terms of interpretability and controllability while maintaining generation quality.
Warnings: This paper contains depictions and discussions of explicit and other inappropriate content that some readers may find disturbing, distressing, and/or offensive.
@@ 1 Introduction
Diffusion models [11, 39], as the current state-of-the-art (SOTA) generative paradigm, drive the development of text-to-image (T2I) generation systems. These T2I models are guided by textual prompts and aim to generate realistic images that align with the provided text descriptions, finding widespread application in fields such as art creation [44] and graphic design [51]. Between 2022 and 2023 alone, T2I models generated over 15 billion images [6]. While T2I models demonstrate significant potential in terms of generation performance, they also face risks of unsafe generation. For instance, the BBC reported that AI-generated child sexual abuse materials were being widely sold over the internet, severely violating ethical and legal norms [5]. Therefore, ensuring that the content generated by T2I models adheres to usage policies [8, 25] is an urgent necessity.
@@
To address this issue, many commercial T2I online services have implemented various safeguards. For instance, DALL-E [30] uses a Moderation API to detect input prompts that may violate usage policies. Stable Diffusion is equipped with a SafetyChecker [31] to filter out generated images containing explicit content. However, as passive defense mechanisms, input and output moderation typically offer coarse-grained protection, which may cause T2I models to overly intercept and refuse to generate any images. In contrast, safe generation methods [17, 7, 15] focus on erasing inappropriate concepts from the model’s internal representations. These methods aim to suppress the emergence of these concepts in the generated content, rather than indiscriminately blocking inputs or outputs. For instance, SafeGen [17] fine-tunes the model to remove visual representations linked to explicit concepts, ensuring that the generated content does not reflect these concepts. Such methods offer more fine-grained protection for T2I models, enabling them to produce high-quality, safe images even when given prompts with inappropriate concepts.
@@
However, despite the fact that existing safe generation methods are capable of suppressing specific target concepts to certain extent, their robustness remains limited [43, 46]. Considering that semantic consistency between the input text and the generated image is a fundamental requirement for T2I models, which means that inappropriate prompts can be the primary source of unsafe generations by T2I models. Existing methods usually focus on fine-tuning the denoiser of a T2I model or modifying the denoising process to erase inappropriate concepts from the visual representation, neglecting the sanitization of inappropriate concepts in the prompt embeddings. This exposes their limitations, as these methods typically only suppress explicitly deleted concepts and lack sufficient generalization ability when handling synonyms or related concepts. Adversarial attacks on T2I models, such as SneakyPrompt [49], exploit this weakness by using synonym substitutions to bypass safeguards, leading to unsafe generation. Therefore, a natural question arises: Could we propose a robust safety generation framework by erasing inappropriate concepts at the source (i.e., prompt embeddings)?
Figure 1: ES operates by sanitizing prompt embeddings to generate safe images and providing interpretability to indicate harmful tokens.
@@
Our work. In this paper, we provide an affirmative answer by introducing a novel, vision-agnostic safe generation framework, Embedding Sanitizer (ES). As shown in Figure 1, ES operates as a plug-and-play module applied after the text encoder, without necessitating modifications to any T2I model components. This design ensures compatibility, allowing ES to function independently or integrate seamlessly with other protective measures. ES focuses on erasing representations of inappropriate concepts from prompt embeddings, producing sanitized embeddings that guide the T2I model in generating appropriate content. A key feature of ES is its internal scoring network, which assigns a score to each token in an input prompt to gauge its potential harmfulness. For example, when processing the prompt ”A person in a birthday suit,” the token ”suit” receives a high score, indicating it as a significant contributor to inappropriate content generation. During the sanitization phase, ES dynamically adjusts the intensity of sanitization based on these scores, minimizing the impact on low-score (harmless) tokens while preserving strong sanitization effects on high-score (harmful) tokens. The scoring mechanism and adaptive sanitization function not only enhance the identification and management of inappropriate content but also provide interpretability and controllability in the generation process.
@@
ES is a plug-and-play deep learning model that accesses only the text encoder during training, without involving other T2I model components. This design supports the use of a customized objective function for efficient end-to-end training. As a vision-agnostic safe generation framework, ES does not rely on image data for training. Instead, training begins with the selection of target concepts for erasure (e.g., nude) and their corresponding anchor concepts (e.g., dressed), providing explicit learning targets for sanitization. To augment training data, we use randomly synthesized prompts to pair toxic prompts $P_{t}$ (e.g., A nude person) with clean prompts $P_{c}$ (e.g., A dressed person). This approach produces abundant, diverse contextual samples, enabling ES to effectively learn to identify and sanitize inappropriate content across various scenarios. During training, a mean squared error loss is employed to encourage ES to align the embedding of the toxic prompt $P_{t}$ with that of the clean prompt $P_{c}$. Post-training, ES sanitizes input prompts in real-time to produce clean embeddings, ensuring that generated content adheres to usage policies.
@@
  • We propose and investigate a vision-agnostic safe generation method, Embedding Sanitizer (ES), designed as a plug-and-play module that sanitizes prompt embeddings in a controllable manner and provides interpretability for inappropriate elements within prompts. This modular design ensures orthogonal compatibility with existing safeguards, allowing ES to function independently or integrate seamlessly with other protective measures.
  • We introduce a synthetic data-driven training strategy that supplies ES with a nearly limitless variety of contextual training data. By specifying target and anchor concepts for removal, we reformulate the sanitization task as a transformation from target to anchor concepts, enabling efficient end-to-end training of ES.
  • We conduct a comprehensive evaluation of ES on five prompt datasets and compare it with nine latest safeguard baselines. Results demonstrate that ES achieves SOTA robustness while maintaining content quality, validating that prompt embedding sanitization can significantly enhance resilience to adversarial attacks. To facilitate future studies, we open-source our code in the following repository: https://anonymous.4open.science/r/Embedding-Sanitizer-166E/.
Ethical Considerations. For inappropriate content displayed in the manuscript, we apply masking measures to ensure that readers are not directly exposed to sensitive materials, thereby adhering to ethical standards for handling such content.
@@ 2.1 Text-to-Image (T2I) Generation Models
Diffusion models have recently become a SOTA generative paradigm [10], employing a probabilistic framework to generate high-fidelity images from noisy representations. The training of diffusion models is divided into the diffusion process and the denoising process. In the diffusion process, the model will gradually add noise to the image so that it eventually approximates the standard Gaussian distribution. In the denoising process, the denoising network predicts and removes the added noise, progressively recovering the original image. The training objective of diffusion models is to minimize the difference between the predicted noise and the true noise at each timestep, typically expressed using a simple mean squared error (MSE) loss:
$\mathcal{L}_{\text{MSE}}=\mathbb{E}_{x,c,\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t,c\right)\right\|_{2}^{2}\right],$ (1)
where $x$ represents the original image, $x_{t}$ is the noisy image at timestep $t$, and $c$ represents the guiding condition, such as a prompt. The variable $\epsilon$ is the randomly sampled Gaussian noise, while $\epsilon_{\theta}$ indicates the predicted noise generated by the denoising network parameterized by $\theta$.
The ability of diffusion models to generate high-fidelity images largely depends on the efficient processing of data in the original high-dimensional pixel space. however, this also results in significant computational overhead. To address this issue, latent diffusion models (LDM)[33] propose a more efficient solution by executing the diffusion and denoising processes in a low-dimensional latent space, significantly reducing computational complexity. Stable Diffusion (SD) [40, 41, 42] is a typical representative of the LDM framework, which mainly consists of three core components:
  • Text Encoder. The text encoder utilizes a pre-trained CLIP model [29] to convert input text prompts into dense semantic embeddings. These embeddings guide the generation process through cross-attention mechanism, ensuring that the generated images align with the semantic meaning of the input text.
  • Image Decoder. The image decoder uses a pre-trained variational autoencoder (VAE) model [14] to map high-dimensional images into a low-dimensional latent space, improving computational efficiency. During the inference stage, the denoised latent representations generated by the diffusion process are decoded by the VAE into high-resolution images.
  • Denoiser. The denoising network uses the U-Net architecture [34], which takes noisy latent representations and text embeddings as input and progressively predicts and removes the noise at each time step, refining the latent representations into coherent image representations.
Compared to earlier generative models, Diffusion models offer a more stable training process, overcoming common issues such as mode collapse [2]. However, despite the breakthroughs in the generation performance of T2I models, challenges remain in terms of safety [52, 50]. Ensuring that the generated images comply with safety guidelines and preventing the production of inappropriate, harmful, or biased content are still pressing issues that need to be addressed.
@@ 2.2 Adversarial Attacks on T2I Models
With the support of large-scale image-text pair datasets such as LAION-5B [37], the performance of T2I models achieves significant progress. However, certain content within these datasets may contain sensitive information (such as adult content or hate symbols), which the models may inadvertently memorize during training, subsequently generating harmful images. Despite the integration of various safeguards in T2I models to avoid this issue, they remain vulnerable to adversarial attacks.
Adversarial attacks against T2I models aim to create adversarial prompts that bypass or evade the model’s safeguards, generating inappropriate images that do not meet expected standards. These attacks typically manipulate or modify the input prompts, introducing adversarial elements while preserving semantic similarity to the original prompts, thereby rendering the model’s safeguards ineffective. Based on the adversary’s knowledge of the target model, adversarial attacks can be categorized into white-box attacks and black-box attacks. In white-box attacks, the adversary has full access to the internal information of the target model, including its architecture, parameters, and gradient calculations. This allows the adversary to directly use the internal knowledge of the model to construct adversarial samples. For example, by utilizing gradient-based optimization techniques, the adversary can identify a pseudo token to compel the model to generate specific inappropriate content [26].
In contrast, black-box attacks are conducted without access to the internal structure or parameters of the model, making them more reflective of real-world threat scenarios. Adversaries typically rely on locally deployed surrogate models to optimize adversarial prompts, leveraging the transferability of prompts to attack the target model. For instance, SneakyPrompt [49] uses a Stable Diffusion model equipped with an output detector as a surrogate model and applies reinforcement learning algorithms to automatically search for substitutes for sensitive words in prompts. These substitutes are optimized using a reward function based on safety compliance and semantic fidelity, encouraging the generated adversarial prompts to bypass safeguards. Ring-A-Bell [43] utilizes a CLIP model to extract context-independent sensitive concept vectors and add them to the target prompts, and then uses a genetic algorithm to search for semantically similar adversarial prompts. Similarly, MMA-Diffusion[46] employs gradient-driven optimization and sensitive word regularization to construct adversarial prompts that contain no sensitive words while maintaining semantic similarity to the target prompt.
@@ 3 Related Work and Motivation
In this section, we review and summarize existing safeguards in the T2I model. We then summarize the properties that an ideal safe generation framework should possess. By analyzing these properties, we highlight the limitations of existing safe generation methods, which motivates us to propose a new defense framework that satisfies these properties.
@@ 3.1 Safeguards for T2I Models
Safe Generation. Safe generation, also known as concept erasure [21], is a type of machine unlearning method [4] applied at the image generation stage. This method aims to erase inappropriate target concepts from the model’s internal representations, ensuring appropriate content output even when users provide malicious prompts. Negative Prompt [1] is a concept suppression mechanism integrated in Stable Diffusion that enables the user or the system to specify concepts that should not be generated in order to avoid generating inappropriate content. In the denoising process, the method suppresses the inappropriate content by introducing the embedding of inappropriate concepts to add noise to the intermediate denoising result. SLD [35] combines classifier-free guided text conditioning with inappropriate concepts to suppress the generation of these concepts similarly to Negative Prompt. ESD [7] fine-tunes the cross-attention or non-cross-attention layers of the denoiser on target concepts, embedding the functionality of Negative Prompt into the model parameters. Similarly, CA[15] fine-tunes either the full denoiser or only its cross-attention on pairs of target prompts and anchor images, aiming to link inappropriate target concepts with safe and appropriate images. Safe-CLIP[27] relearns a safe CLIP model through contrastive learning between text and image pairs, aiming to break associations between inappropriate textual and visual concepts. SafeGen[17] focuses on eliminating sexually explicit content by fine-tuning the self-attention layers of the denoiser in a text-independent manner, forcing the model to apply a mosaic effect to any generated explicit content.
@@
Properties Analysis of Safe Generation. To comprehensively guard against potential evasion strategies, such as adversarial attacks, the safe generation framework for T2I models should meet the following core criteria [20]: (i) Effectiveness. The framework must effectively erase target concepts so that they do not appear in the generated image. (ii) Robustness. The framework should possess strong robustness, capable of suppressing synonymous concepts of the target and maintaining its effectiveness against adversarial attacks. (iii) Specificity. When generating non-target concepts, the output images should align semantically with the text descriptions, while maintaining a distribution consistent with the original model. (iv) Fidelity. After erasing target concepts, the quality of the generated images should not be significantly impacted, ensuring that the resulting images remain high-quality and usable.
@@
Furthermore, to enhance defense performance and flexibility, an ideal safe generation framework should satisfy the following advanced criteria: (v) Interpretability. The framework should be able to reveal the contributions of tokens to the generation of inappropriate content, enabling T2I model providers to understand the extent of specific inputs’ impact on outputs, facilitating debugging and improvement. (vi) Compatibility. The framework should operate independently of existing defense systems, allowing for seamless integration and collaboration with various protective measures. (vii) Controllability. The framework should allow for control over the degree of erasure of target concepts during use, enabling T2I model providers to dynamically adjust the defense strength based on specific application needs.
3.2 Limitations of Existing Safeguards.
Despite significant progress in the safeguards of T2I models, existing mechanisms still exhibit notable limitations when dealing with complex adversarial attacks [13]. First, input moderation effectively intercepts prompts containing common sensitive terms but performs inadequately when handling synonyms, paraphrases, and adversarial prompts [47]. Second, output moderation is constrained by the detection scope and performance of content detectors [28], which may fail to detect attacks such as prompt dilution [43]. Furthermore, input and output moderation typically provide only coarse-grained protection, often resulting in the blind interception of inappropriate content, which causes T2I models to refuse to generate any content, negatively impacting user experience [45, 16].
In contrast, safe generation methods provide finer-grained control over the target concepts, allowing for the generation of safe and appropriate outputs even in the presence of prompts containing inappropriate concepts. This is because safe generation methods focus on eliminating or suppressing the generation of inappropriate concepts rather than completely blocking inputs or outputs, ensuring that users receive meaningful and compliant results when using the model. Although safe generation methods perform well in actively suppressing the generation of inappropriate concepts, they still lack robustness and are vulnerable to various adversarial prompts. Additionally, existing methods face further challenges in terms of interpretability, composability, and multi-concept erasure. Table I summarizes and compares the characteristics of influential existing safe generation methods. In terms of interpretability, existing methods primarily focus on suppressing the generation of specific concepts, making it difficult to pinpoint which tokens in the prompt trigger inappropriate outputs. This lack of interpretability hampers T2I service providers’ ability to effectively adjust and optimize defensive measures. Regarding compatibility, while SafeGen[17] and Safe-CLIP[27] claim seamless integration with other safeguards, SafeGen may encounter objective function conflicts with other fine-tuning-based methods such as ESD and AC. The secure text encoder built by Safe-CLIP can weaken the effectiveness of safe generation methods relying on embedding inappropriate concepts, such as Negative Prompt and SLD. In terms of controllability, existing methods often fail to dynamically adjust the defense intensity based on real-world application needs, making it difficult to balance defense performance and generation quality, which in turn limits the quality and diversity of generated outputs. These limitations motivate us to develop more robust safe generation frameworks to counteract the threats posed by increasingly sophisticated adversarial attacks.
@@
NP
SLD
ESD U-Net
CA U-Net
Safe-CLIP Text-Encoder
SafeGen U-Net

### v2

@@ Abstract
Given that semantic consistency between input text and output image is a core requirement of T2I models, we identify that textual representations are likely the primary source of unsafe generation. To this end, we propose Embedding Sanitizer (ES), which enhances the safety of T2I models by sanitizing inappropriate concepts in prompt embeddings. To our knowledge, ES is the first interpretable safe generation framework that assigns a score to each token in the prompt to indicate its potential harmfulness. In addition, ES adopts a plug-and-play modular design, offering compatibility for seamless integration with various T2I models and other safeguards. Evaluations on five prompt benchmarks show that ES outperforms eleven existing safeguard baselines, achieving state-of-the-art robustness while maintaining high-quality image generation.
@@ 1 Introduction
Diffusion models [15, 42], as the current state-of-the-art (SOTA) generative paradigm, drive the development of text-to-image (T2I) generation systems. Between 2022 and 2023 alone, the T2I models generated more than 15 billion images [7, 3]. Although these models exhibit remarkable generative capabilities, they also pose significant risks of unsafe content generation. For example, BBC reported that AI-generated child sexual abuse materials were widely distributed online, constituting severe violations of ethical and legal standards [6]. Therefore, it is an urgent necessity to ensure that the content generated by the T2I models adheres to usage policies [10].
@@
To address this issue, many commercial T2I online services have implemented various safeguards. For instance, DALL-E [33] uses a Moderation API to detect input prompts that may violate usage policies. Stable Diffusion [43] is equipped with a SafetyChecker [34] to filter out generated images containing explicit content. However, as passive defense mechanisms, input and output moderation typically offer coarse-grained protection, which may cause T2I models to overly intercept and refuse to generate any images. In contrast, safe generation methods [21, 8, 19] focus on erasing inappropriate concepts from the model’s internal representations. These methods aim to suppress the emergence of these concepts in the generated content, rather than indiscriminately blocking inputs or outputs. Such methods offer more fine-grained protection for T2I models, enabling them to produce safe images even when given prompts with inappropriate concepts.
Figure 1: ES sanitizes prompt embeddings to ensure safe image generation while also identifying inappropriate tokens.
@@
However, despite the fact that existing safe generation methods can suppress the target concept to some extent, their robustness remains limited [47, 50]. As shown in Figure 1, existing methods typically focus on fine-tuning the denoiser or modifying the denoising process to erase inappropriate concepts from the visual representation, while neglecting the sanitization of inappropriate concepts in the text representation. This leads to their ability to typically suppress explicitly erased concepts but lack sufficient generalization when handling synonyms or related concepts. Adversarial attacks, such as SneakyPrompt [52], exploit this weakness by using synonym replacements or subtle expressions to bypass the safeguards, resulting in unsafe generation. Given that semantic consistency between input text and generated image is a core requirement for T2I models, we find that text representation may be the primary source of unsafe generation. Therefore, a natural question arises: Can we propose a robust safe generation framework by erasing inappropriate concepts from the source?
Our work. In this paper, we provide an affirmative answer by introducing a novel, vision-agnostic safe generation framework, Embedding Sanitizer (ES). Our method focuses on sanitizing text representations in prompt embedding space rather than raw text space for two main reasons. First, the prompt embedding, as the direct input guiding image generation, contains richer semantic information compared to the discrete raw text, and its sanitization provides direct and effective guidance for safe generation. Second, the inherent relational structure in the embedding space ensures that the distance between synonymous (e.g., ”nude” and ”naked”) is typically smaller [16, 52], thus offering better generalization for ES when handling tokens not seen during training to ensure high coverage of inappropriate concepts.
@@
As shown in Figure 1, ES operates as a plug-and-play module applied after the text encoder, requiring no modifications to any T2I model components. This design ensures compatibility, enabling ES to function independently or seamlessly integrate with other safeguards. A key function of ES is its internal scoring network, which assigns a score to each token in an input prompt to gauge its potential harmfulness. For example, when processing the prompt ”A person in a birthday suit,” the token ”suit” receives a high score, indicating it as a significant contributor to inappropriate content generation and needs to be sanitized. In contrast, other tokens receive low scores, indicating that they are benign tokens and should be retained. During the sanitization phase, ES dynamically adjusts the intensity of sanitization based on these scores to minimize the impact on low-score (harmless) tokens, while ensuring that strong sanitization effects on high-score (harmful) tokens. The scoring mechanism and dynamic sanitization function not only enhance the identification and management of inappropriate content but also provide interpretability in the generation process.
@@
ES is a deep learning model that accesses only the text encoder during training, without involving other T2I model components. This design supports the use of a customized objective function for efficient end-to-end training. As a vision-agnostic safe generation framework, ES does not rely on image data for training. Instead, training begins with the selection of target tokens for erasure (e.g., nude) and their corresponding anchor tokens (e.g., dressed), providing explicit learning targets for embedding sanitization. To augment training data, we use randomly synthesized prompts to pair toxic prompts $P_{t}$ (e.g., ”A nude person”) with anchor prompts $P_{a}$ (e.g., ”A dressed person”), allowing for the construction of virtually unlimited training samples. This approach ensures a diverse and contextually rich dataset, enabling ES to effectively learn to detect and sanitize inappropriate content across various textual inputs. We employ a mean squared error (MSE) loss to encourage ES to align the embeddings of $P_{t}$ with those of $P_{a}$ to ensure that the generated content complies with usage policies.
@@
$\bullet$ We propose the first interpretable safe generation framework, Embedding Sanitizer (ES), which quantifies the harmfulness of each token in the prompt and sanitizes the prompt embedding in a controllable manner. The modular design of ES ensures compatibility, allowing seamless integration with various T2I models and other safeguard mechanisms.
$\bullet$ We introduce a synthetic data-driven training strategy that provides ES with virtually unlimited diverse contextual training samples. By specifying target tokens to be removed and anchor tokens for guidance, we reformulate the sanitization task as a semantic transformation from the target token to the anchor token, enabling efficient end-to-end training of ES.
$\bullet$ We conduct a comprehensive evaluation of ES on five prompt datasets and compare it against eleven existing safeguard baselines. Results show that ES achieves state-of-the-art robustness while maintaining generation quality, demonstrating that prompt embedding sanitization significantly enhances the safety of T2I generation. We open-source our code in the following repository: https://anonymous.4open.science/r/Embedding-Sanitizer-166E/.
@@ 2.1 Text-to-Image (T2I) Generation Models
Diffusion models have recently become a SOTA generative paradigm [14], employing a probabilistic framework to generate high-fidelity images from noisy representations. The training of diffusion models is divided into the diffusion process and the denoising process. In the diffusion process, the model will gradually add noise to the image so that it eventually approximates the standard Gaussian distribution. In the denoising process, the denoising network predicts and removes the added noise, progressively recovering the original image. The training objective of diffusion models is to minimize the difference between the predicted noise and the true noise at each timestep. Latent diffusion models (LDM) [36] propose a more efficient solution by executing the diffusion and denoising processes in a low-dimensional latent space, significantly reducing computational complexity. Stable Diffusion (SD) [43, 44, 45] is a typical representative of the LDM framework, which mainly consists of three core components:
$\bullet$ Text Encoder. The text encoder utilizes a pre-trained CLIP model [32] to convert input text prompts into embeddings. These embeddings guide the generation process through cross-attention mechanism, ensuring that the generated images align with the semantic meaning of the input text.
$\bullet$ Image Decoder. The image decoder uses a pre-trained variational autoencoder (VAE) model [18] to map high-dimensional images into a low-dimensional latent space, improving computational efficiency. During the inference stage, the denoised latent generated by the diffusion process are decoded by the VAE into images.
$\bullet$ Denoiser. The denoising network uses the U-Net architecture [37], which takes noisy latent and text embeddings as input and progressively predicts and removes the noise at each time step, refining the latent into coherent image representations.
@@ 2.2 Adversarial Attacks on T2I Models
With the support of large-scale image-text pair datasets such as LAION-5B [40], the performance of T2I models achieves significant progress. However, certain content within these datasets may contain sensitive information (such as adult content or hate symbols), which the models may inadvertently memorize during training, subsequently generating harmful images [4]. Despite the integration of various safeguards in T2I models to avoid this issue, they remain vulnerable to adversarial attacks.
Adversarial attacks against T2I models aim to create adversarial prompts that bypass or evade the model’s safeguards, generating inappropriate images that do not meet expected standards. These attacks typically manipulate or modify the input prompts, introducing adversarial elements while preserving semantic similarity to the original prompts, thereby rendering the model’s safeguards ineffective. Based on the adversary’s knowledge, adversarial attacks can be categorized into white-box attacks and black-box attacks. In white-box attacks, the adversary has full access to the internal information of the target model, including its architecture, parameters, and gradient calculations. This allows the adversary to directly use the internal knowledge of the model to construct adversarial samples. For example, by utilizing gradient-based optimization techniques, the adversary can identify a pseudo token to compel the model to generate specific inappropriate content [29].
In contrast, black-box attacks are conducted without access to the internal information of the model, which makes them more reflective of real-world threat scenarios. Adversaries typically rely on locally deployed surrogate models to optimize adversarial prompts, leveraging the transferability of prompts to attack the target model. For instance, SneakyPrompt [52] uses a Stable Diffusion as a surrogate model and applies reinforcement learning algorithms to automatically search for substitutes for sensitive tokens in the prompt. These substitutes are optimized using a reward function based on safety compliance and semantic fidelity, encouraging the generated adversarial prompts to bypass safeguards. Ring-A-Bell [47] utilizes a CLIP model to extract context-independent sensitive concept vectors and add them to the target prompts, and then uses a genetic algorithm to search for semantically similar adversarial prompts. Similarly, MMA-Diffusion[50] employs gradient-driven optimization and sensitive token regularization to construct adversarial prompts that contain no sensitive words while maintaining semantic similarity to the target prompt.
@@ 3.1 Safeguards for T2I Models
Safe Generation. Safe generation, also known as concept erasure [25], is a type of machine unlearning method [5] applied at the image generation stage. This method aims to erase inappropriate target concepts from the model’s internal representations, ensuring appropriate content output even when users provide malicious prompts. Negative Prompt [1] is a concept suppression mechanism integrated in Stable Diffusion that enables the user or the system to specify concepts that should not be generated in order to avoid generating inappropriate content. In the denoising process, the method suppresses the inappropriate content by introducing the embedding of inappropriate concepts to add noise to the intermediate denoising result. SLD [38] combines classifier-free guided text conditioning with inappropriate concepts to suppress the generation of these concepts similarly to Negative Prompt. ESD [8] fine-tunes the cross-attention or non-cross-attention layers of the denoiser on target concepts, embedding the functionality of Negative Prompt into the model parameters. POSI [49] focuses on fine-tuning a large language model to play the role of a prompt optimizer, rewriting toxic prompts into clean prompts in raw text space rather than embedded space. Safe-CLIP [30] relearns a safe CLIP model through contrastive learning between text and image pairs, aiming to break associations between inappropriate textual and visual concepts. SafeGen [21] focuses on eliminating sexually explicit content by fine-tuning the self-attention layers of the denoiser in a text-independent manner, forcing the model to apply a mosaic effect to any generated explicit content. RR [46] adopts the idea of backdoor attacks, treating harmful tokens as triggers and mapping their semantics to alternative prompts that remove the toxic content. Moderator [48] formulates the modeling of harmful concepts as a task vector at the parameter level, and obtains a safer T2I model by subtracting this vector from the original model.
3.2 Properties Analysis of Safe Generation
@@
To comprehensively guard against potential evasion strategies, such as adversarial attacks, the safe generation framework should meet the following core criteria [24]: (i) Effectiveness. The framework effectively erases target concepts to ensure that they do not appear in the generated image. (ii) Robustness. The framework should be able to suppress synonymous concepts of the target and maintain its effectiveness against adversarial attacks. (iii) Specificity. When generating benign (non-target) concepts, the output images should semantically align with the corresponding text descriptions. (iv) Fidelity. The quality of benign image generation should not be significantly degraded after erasing the target concepts.
@@
Furthermore, to enhance defense performance and flexibility, an ideal safe generation framework should satisfy the following advanced criteria: (v) Interpretability. The framework should be able to reveal the contributions of tokens to the generation of inappropriate content, helping T2I model providers understand their impact and improve the system. (vi) Compatibility. The framework should work independently of existing safeguards, allowing easy integration with different protection measures. (vii) Controllability. The framework should allow the provider to control the level of erasure of target concepts, enabling flexible defense levels based on application needs. For example, for malicious users with a high number of violations, a higher erasure strength can be set to enhance content security management.
@@
NP [1]
SLD [38]
ESD [8] U-Net
POSI [49]
Safe-CLIP [30] Text-Encoder
SafeGen [21] U-Net
RR [46] Text-Encoder
Moderator [48] U-Net
@@
3.3 Limitations of Existing Safeguards.
Despite significant progress in the safeguards of T2I models, existing mechanisms still exhibit notable limitations when dealing with complex adversarial attacks [17]. First, input moderation effectively intercepts prompts containing common sensitive terms but performs poorly when handling synonyms, paraphrases, and adversarial prompts [51]. Second, output moderation is constrained by the detection scope and performance of content detectors [31], which may fail to detect attacks such as prompt dilution [47]. Furthermore, input and output moderation typically provide only coarse-grained protection, often resulting in the blind interception of inappropriate content, which causes T2I models to refuse to generate any content, negatively impacting user experience [49, 20].
In contrast, safe generation methods provide finer-grained control over the target concepts, allowing for the generation of safe and appropriate outputs even in the presence of prompts containing inappropriate concepts. This is because safe generation methods focus on eliminating or suppressing the generation of inappropriate concepts rather than completely blocking inputs or outputs, ensuring that users receive meaningful and compliant results when using the model. Although safe generation methods perform well in actively suppressing the generation of inappropriate concepts, they still lack robustness and are vulnerable to various adversarial prompts. Additionally, existing methods face further challenges in terms of interpretability, composability, and Controllability. Table I summarizes and compares the characteristics of influential existing safe generation methods. In terms of interpretability, existing methods tend to blindly suppress the generation of target concepts, making it difficult to pinpoint which tokens in the prompt trigger inappropriate outputs. This lack of interpretability hampers T2I service providers’ ability to effectively adjust and optimize defensive measures. Regarding compatibility, while SafeGen[21] and Safe-CLIP[30] claim seamless integration with other safeguards, SafeGen may encounter objective function conflicts with other fine-tuning-based methods such as ESD. The secure text encoder built by Safe-CLIP can weaken the effectiveness of safe generation methods relying on embedding inappropriate concepts, such as Negative Prompt and SLD. In terms of controllability, existing methods often fail to dynamically adjust the defense intensity based on real-world application needs, making it difficult to balance defense performance and generation quality, which in turn limits the quality and diversity of generated outputs. These limitations motivate us to develop more robust safe generation frameworks to counteract the threats posed by increasingly sophisticated adversarial attacks.